Axolotl:

Please help me! I need help with Trigonometry: Plane A takes off at a 13° angle from the runway, and plane B takes off at a 9° angle from the runway. Which plane reaches a greater horizontal distance from the airport when the plane reaches an altitude of 10,000 feet? Round the solutions to the nearest whole number. Plane A because it was 43,315 feet away Plane A because it was 63,925 feet away Plane B because it was 44,454 feet away Plane B because it was 63,138 feet away

Axolotl:

@jhonyy9 please help

surjithayer:

 \[\frac{ 10000 }{ x1 }=\tan 13\] \[x1=10000 ~\cot~13=10000 \cot(90-77)=10000 \tan77\approx 43315~ft\] \[\frac{ 10000 }{ x2 }=\tan 9\] \[x2=10000 \cot 9=10000\cot(90-81)=10000 \tan 81\approx 63138 ft\] Plane B because it was 63138 ft away.